OpenStudy (anonymous):

Factor 2x^4-8x^3-10x^2 completely

OpenStudy (pfenn1):

\[2x^4-8x^3-10x^2=2x^2(x^2-4x-5)\]Factor \[x^2-4x-5=(x+a)(x+b)\]where a+b=-4 and ab=-5